# Liquidity Fragmentation or Optimization? Analyzing Automated Market Makers Across Ethereum and Rollups

### v2

@@ 4 Empirical Analysis
Figure 4: Overview of WHET-USDC liquidity pools on Uniswap (v3) across Ethereum and its rollups: trading volume, TVL in the current tick and LP returns. The last chart illustrates the reallocation of liquidity required to attain equilibrium. Liquidity that is not allocated to any pool is directed towards staking.
Data.
We collect on-chain data for every swap in the WETH-USDC pools on Uniswap v3, covering the period from 2023 to the first half of 2024. The analysis includes Ethereum and its EVM-compatible L2s, specifically optimistic rollups such as Arbitrum, Optimism, and Base, as well as a ZK-rollups — ZKsync. For Arbitrum, we analyze pools with native and bridged USDC against WETH. LP fee in these pools is 5bps, with the exception of ZKsync pool with 20bps. The trading volume in these pools is shown in Figure 4(a).
Methodology.
For each day and each pool, we find the last swap and, based on the liquidity and current tick, we calculate TVL in the current tick, shown in Figure 4(b). Then, we calculate the returns of the AMM pool based on the TVL in the current tick using Equation (3) and the optimal allocation using Equation (5). We also assume that the LP provides liquidity in ticks around 12% of the current spot price and, based on the finding of Adams [1], the liquidity concentration is 75% higher on L2s compared to Ethereum. The annualized return of such LP position is depicted in Figure 4(c).
Empirical Calibration.
To model the relationship between trading volume and TVL in liquidity pools, we use the function defined in Equation (6). To estimate the parameters $k$ and $\epsilon_{v}$, we apply a logarithmic transformation to linearize the equation:
@@
which can be expressed as a linear regression model
$y=a+bx$ (11)
where $y=\log(\text{Vol})$, $x=\log(\text{TVL})$, $a=\log(k)$, and $b=\epsilon_{v}$.
We perform a linear regression analysis on historical on-chain data for each liquidity pool. The slope of the regression line provides an estimate of $\epsilon_{v}$, while the intercept gives $\log(k)$, from which $k$ can be obtained by exponentiating the intercept: $k=e^{a}$. The elasticity parameter $\epsilon_{v}$ indicates how sensitive trading volume is to changes in TVL. An elasticity greater than 1 ($\epsilon_{v}>1$) suggests a highly elastic volume, where trading activity increases significantly as liquidity grows. Conversely, an elasticity of 1 or less ($\epsilon_{v}\leq 1$) indicates inelastic volume, where trading volume grows slower than TVL.
Chain Elasticity ($\epsilon_{v}$) Scaling constant ($k$)
Ethereum -0.12079669 3053256390.85974741
Arbitrum -0.17739562 3945020961.93421507
Arbitrum2 -0.14315817 919336413.99392521
Base 1.04506956 0.11842139
Optimism -0.17837105 236925148.26732501
ZKsync 0.65420951 4.70075502
Table 1: Elasticity and scaling constants for Ethereum and selected rollups.
@@ 4.1 Interpretation of Elasticity and Scaling Constant Results
Table 1 presents the estimated elasticity of volume with respect to virtual TVL and the scaling constant for six different chains. These metrics are derived from a log-log regression model that captures the relationship between TVL and trading volume. The key insights from the results are discussed below:
The elasticity of volume with respect to TVL for Ethereum is estimated to be $-0.12$. This suggests that, contrary to expectations, an increase in TVL on Ethereum is associated with a slight decrease in trading volume. However, the scaling constant ($k=3.05\times 10^{9}$) indicates a high baseline volume relative to TVL, suggesting that Ethereum has a significant trading volume regardless of TVL changes.
Both Arbitrum and Arbitrum (e) pools exhibit negative elasticity values of $-0.18$ and $-0.14$, respectively. These results imply that an increase in TVL leads to a decrease in trading volume on these chains. The scaling constants are $3.95\times 10^{9}$ and $9.19\times 10^{8}$, respectively, indicating that Arbitrum has a higher baseline volume compared to Arbitrum (e). Optimism displays a negative elasticity of $-0.18$, similar to Arbitrum. The scaling constant for Optimism is $2.37\times 10^{8}$, indicating a lower baseline volume compared to Arbitrum and Ethereum.
@@
The elasticity results highlight the variability in the relationship between TVL and trading volume across different chains. While Base and ZKsync — new blockchains (launched in 2023) with new deployments of Uniswap (v3) show positive elasticities, suggesting that TVL growth drives volume, Ethereum, Arbitrum, and Optimism — well established and older chains - show negative elasticities, which may be attributed to differences in user behavior, liquidity distribution, or (lack of) protocol incentives on these chains. Both Arbitrum and Optimism were launched in 2021, and Ethereum in 2015.
@@ 4.2 Interpretation of Optimal Allocation Results
Table 2 presents the calculated results for the new LP willing to allocate her liquidity among the pools on Ethereum and its L2s on 30th April 2024, assuming 3.47% staking rate of ETH. The first observation is that the AMM pool on Ethereum is oversubscribed with a return rate lower than the staking rate. Thus, the current LPs would yield higher returns by reallocating their capital to L2s, or to staking. Second, the highest return presents the pool on Optimism — 22.12%, however, given the trading volume on Optimism, the highest capital allocation in optimal LP strategy is achieved for the pools on Arbitrum and BASE, approx. 80k USD. The returns of LP after allocating capital to the pools on L2s are presented in the last column of Table 2 and ranges between 6 and 8.5%, significantly lower compared to the current returns of 12 to 22%.
@@ 5 Discussion
The natural question that arises from our analysis is why LPs do not reallocate their capital to AMMs on L2s, or staking. The reasons cannot be costs of the token transfers between chains or gas fees, as the observed differences in the returns exceed these costs. As all studied pools have the same market risks and operate on the same AMM, the LPs’ reluctance toward L2s must stem from the security concerns such as centralized sequencers of rollups, L2 upgradeability, or other risks of using this new technology. Similar risk-related preferences were observed among AMM traders on L2s [7].
As L2s are cheaper and faster chains compared to Ethereum, LPs can adjust their CLMM positions more often, resulting in higher liquidity concentration on L2s [1], and higher returns from liquidity provisions. The LP rewards should not only compensate for the missed opportunity not to stake the capital (loss-versus-staking) [11] and for impermanent loss. The AMM pools with more volatile tokens should yield higher rewards over staking than less volatile cryptocurrency pairs.
@@ 6 Conclusions
Using Lagrangian optimization, we developed a model for the optimal capital allocation to staking and liquidity provisions across AMMs on Ethereum and its L2s. We showed that the returns of the AMM liquidity provision converge to the staking rate. In addition, we modeled and measured the elasticity of trading volume with respect to TVL at the WETH-ETH pools at Uniswap (v3) across Ethereum and rollups. Our finding indicate that at the well established blockchains - Ethereum, Arbitrum and Optimism, an increase in TVL is not associated with an increase in trading volume. In contrast, emerging blockchains, Base and ZKsync, exhibit a positive elasticity value, indicating that the volume of trade is positively correlated with TVL on these chains.
Lastly, we empirically compared profitability of liquidity provisions to WETH-USDC pools, and observed that the Ethereum pool, compared to corresponding pools on L2s, is oversubscribed and often yields lower returns than staking Ether. Using historical trading volumes and TVLs, we calculated the optimal liquidity allocation for the new LP as well as the optimal capital reallocation across Ethereum and L2s. Specifically, we found that over 66% of Ethereum liquidity should be reallocated to rollups in order to maximize LP returns and to attain equilibrium with staking rates.
Acknowledgements

### v3

@@ 4 Empirical Analysis
Figure 4: Overview of WETH-USDC liquidity pools on Uniswap (v3) across Ethereum and its rollups: trading volume, TVL in the current tick and LP returns. The last chart illustrates the reallocation of liquidity required to attain equilibrium. Liquidity that is not allocated to any pool is directed towards staking.
Data.  For our empirical analysis, we focus on the WETH-USDC pools on Uniswap v3, as they exhibit the highest trading volume and TVL. Additionally, we always examine the WETH-USDC pool on Ethereum and across each L2 to ensure that all LPs are exposed to the same market risks. We collect on-chain data for every swap in these pools, covering the period from 2023 to the first half of 2024. The analysis includes Ethereum and its EVM-compatible L2s, specifically optimistic rollups such as Arbitrum, Optimism, and Base, as well as a ZK-rollups — ZKsync. For Arbitrum, we analyze pools with native USDC and bridged USDC.e against WETH. The Arbitrum pool with bridged USDC.e we further denote as Arbitrum (e). LP fee in these pools is 5bps, with the exception of ZKsync pool with 20bps. The trading volume in these pools is shown in Figure 4(a).
Methodology.  For each day and each pool, we find the last swap and, based on the liquidity and current tick, we calculate TVL in the current tick, shown in Figure 4(b). Then, we calculate the returns of the AMM pool based on the TVL in the current tick using Equation (3) and the optimal allocation using Equation (5). We also assume that the LP provides liquidity in ticks around 12% of the current spot price and, based on the finding of Adams [2], the liquidity concentration is 75% higher on L2s compared to Ethereum. The annualized return of such LP position is depicted in Figure 4(c).
Empirical Calibration.  To model the relationship between trading volume and TVL in liquidity pools, we use the function defined in Equation (6). To estimate the parameters $k$ and $\epsilon_{v}$, we apply a logarithmic transformation to linearize the equation:
@@
which can be expressed as a linear regression model $y=a+bx$ where $y=\log(\text{Vol})$, $x=\log(\text{TVL})$, $a=\log(k)$, and $b=\epsilon_{v}$.
Chain Elasticity ($\epsilon_{v}$) Scaling constant ($k$) $R^{2}$ SE($\epsilon_{v}$) SE($k$)
Ethereum -0.121 3.05e+09 0.012 0.101 2.091
Arbitrum -0.177 3.95e+09 0.039 0.118 2.348
Arbitrum (e) -0.143 9.19e+08 0.010 0.128 2.488
Base 1.045 0.118 0.895 0.033 0.539
Optimism -0.178 2.37e+08 0.026 0.100 1.757
ZKsync 0.654 4.701 0.556 0.054 0.746
Table 1: Elasticity analysis across Ethereum and selected rollups. $\epsilon_{v}$ represents the elasticity coefficient, $k$ the scaling constant, $R^{2}$ the goodness-of-fit, and SE denotes standard errors for $\epsilon_{v}$ and $k$.
@@ 4.1 Interpretation of Elasticity and Scaling Constant Results
Table 1 presents the estimated elasticity of volume with respect to virtual TVL and the scaling constant for six different chains. These metrics are derived from a log-log regression model that captures the relationship between TVL and trading volume. The key insights from the results are discussed below.
The elasticity of volume with respect to TVL for Ethereum is estimated to be $-0.12$. This suggests that, contrary to expectations, an increase in TVL on Ethereum is associated with a slight decrease in trading volume. However, the scaling constant indicates a high baseline volume relative to TVL, suggesting that Ethereum has a significant trading volume regardless of TVL changes. Both Arbitrum and Arbitrum (e) pools exhibit negative elasticity values. These results imply that an increase in TVL leads to a decrease in trading volume on these chains. The scaling constants indicate that Arbitrum has a higher baseline volume compared to Arbitrum (e). Optimism displays a negative elasticity, similar to Arbitrum. The scaling constant for Optimism indicates a lower baseline volume compared to Arbitrum and Ethereum. Yet, the $R^{2}$ values are for these chains remain low.
@@
The elasticity results highlight the variability in the relationship between TVL and trading volume across different chains. While Base and ZKsync — new blockchains (launched in 2023) with new deployments of Uniswap (v3) show positive elasticities, suggesting that TVL growth drives volume, Ethereum, Arbitrum, and Optimism — well established and older chains - show negative elasticities with low $R^{2}$ values indicating that the relation between TVL and Volume cannot be easily established. This can be attributed to differences in user behavior, liquidity distribution, or (lack of) protocol incentives on these chains. Arbitrum and Optimism were launched in 2021, and Ethereum in 2015.
@@ 4.2 Interpretation of Optimal Allocation Results
Table 2 presents the calculated results for the new LP willing to allocate \$212k liquidity among the pools on Ethereum and its L2s on 30th April 2024, assuming 3.47% staking rate of ETH. If the new LP would possess more liquidity, the additional liquidity should be allocated to staking. If the LP would possess less liquidity, it should be allocated to the pools with the highest rewards first. The first observation is that the AMM pool on Ethereum is oversubscribed with a return rate lower than the staking rate. Thus, the current LPs would yield higher returns by reallocating their capital to L2s, or to staking. Second, the highest return presents the pool on Optimism — 22.12%, however, given the trading volume on Optimism, the highest capital allocation in optimal LP strategy is achieved for the pools on Arbitrum and BASE, approx. 80k USD. The returns of LP after allocating capital to the pools on L2s are presented in the last column of Table 2 and ranges between 6 and 8.5%, significantly lower compared to the current returns of 12 to 22%.
@@ 5 Discussion
A key question arising from our analysis is why LPs do not reallocate their capital to AMMs on L2s or to staking. Since our methodology ensures that all selected pools share the same market risk exposure, LPs’ reluctance toward L2s likely stems from security concerns, inertia, or a preference for Ethereum’s infrastructure. Potential security concerns include risks associated with centralized sequencers, the upgradeability of L2 protocols, and bridging vulnerabilities. Similar risk-related preferences have been observed among AMM traders on L2s [8].
Cost of Bridging.  The cost of transferring tokens between Ethereum and L2s is minimal—less than 1 basis point (0.01%) of the trade volume when intent-based interoperability protocols [9], such as Across Protocol [1], are used. For instance, in the case of a \$100k transfer (as illustrated in Figure 5), the cost would not exceed \$5 per rollup, with transaction times ranging from just a few seconds to a maximum of 15 minutes [1]. Furthermore, gas fees on Ethereum and L2s do not pose a significant barrier to token transfers, as these costs remain low—below \$2 on Ethereum and approximately \$0.10 on L2s [10].
Time of Bridging.  The time of transfer during which the tokens cannot earn LP fees lasts only seconds (up to 15 minutes), as bridges and intent-based interoperability protocols absorb the risk of L2 transaction not reaching the full finality on Ethereum (hard finality), but assume trust in the centralized sequencers and execute the transfers once the transaction in included L2 block by the sequencer (soft finality[39]. Especially for optimistic rollups this approach reduces the finality time from 7 days of hard finality to 0.2-2.0s of soft finality.
Future Research.  LP rewards should not only compensate for impermanent loss and the opportunity cost of not staking capital (loss-versus-staking) [14] but should also reflect the volatility of the traded assets. AMM pools with more volatile tokens should yield higher returns relative to staking than pools with less volatile pairs. This study focuses on the highly liquid WETH-USDC pool on Uniswap v3. Future research could expand the analysis to other pools, particularly those with lower liquidity or higher volatility (e.g., memecoin pools), where loss-versus-rebalancing (LVR) [26] and impermanent loss may significantly influence LP allocation decisions. A sensitivity analysis that incorporates variations in key parameters, such as TVL, trading volume, and stake rates, or an analysis of other relationships between TVL and trading volume, could provide further insight into the behavior of LP.
@@ 6 Conclusions
The question of whether L2 blockchains cause the liquidity fragmentation of Ethereum has been the subject of extensive debate. Our research indicates that such fragmentation is currently not occurring, but, it could develop in the future—particularly if LPs become aware of potentially higher returns available on L2s. Using Lagrangian optimization, we developed a model for the optimal liquidity provisions across AMMs on Ethereum and its L2s with staking as a reference rate. We showed that the returns of the AMM liquidity provision converge to the staking rate.
In addition, we modeled and measured the elasticity of trading volume with respect to TVL at the WETH-ETH pools at Uniswap (v3) across Ethereum and rollups. Our finding indicate that at the well established blockchains - Ethereum, Arbitrum and Optimism, an increase in TVL is not associated with an increase in trading volume. In contrast, emerging blockchains, Base and ZKsync, exhibit a positive elasticity value, indicating that the volume of trade is positively correlated with TVL on these chains.
Finally, we empirically compared profitability of liquidity provisions to WETH-USDC pools, and observed that the Ethereum pool, compared to corresponding pools on L2s, is oversubscribed and often yields lower returns than staking Ether. Using historical trading volumes and TVLs, we calculated the optimal liquidity allocation for the new LP as well as the optimal capital reallocation across Ethereum and L2s. Specifically, we found that more than 66% of the Ethereum liquidity could be reallocated to rollups to maximize LP returns and to achieve equilibrium with staking rates.
Acknowledgments
@@
Appendix 0.A Proof of Proposition 3
The Lagrangian function is given by:
$\mathcal{L}(\mathbf{w},\lambda)=r_{s}w_{0}+\sum_{i=1}^{n}fk_{i}(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-1}w_{i}+\lambda\left(W-\sum_{i=0}^{n}w_{i}\right).$
Derivative w.r.t $w_{0}$:
$\frac{\partial\mathcal{L}}{\partial w_{0}}=r_{s}-\lambda=0.$
Solving for $\lambda$:
$\lambda=r_{s}.$
For $i=1,\dots,n$:
$\frac{\partial\mathcal{L}}{\partial w_{i}}=fk_{i}\left[(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-1}+(\epsilon_{v}-1)(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-2}w_{i}\right]-\lambda=0.$
Substituting $\lambda=r_{s}$:
$fk_{i}\left[(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-1}+(\epsilon_{v}-1)(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-2}w_{i}\right]=r_{s}.$
Factor out $(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-2}$:
$fk_{i}(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-2}\left[(\text{TVL}_{i}+w_{i})+(\epsilon_{v}-1)w_{i}\right]=r_{s}.$
Since
$(\text{TVL}_{i}+w_{i})+(\epsilon_{v}-1)w_{i}=\text{TVL}_{i}+\epsilon_{v}w_{i},$
we rewrite the equation as:
$fk_{i}(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-2}(\text{TVL}_{i}+\epsilon_{v}w_{i})=r_{s}.$
Dividing both sides by $fk_{i}(\text{TVL}_{i}+\epsilon_{v}w_{i})$, we get:
$(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-2}=\frac{r_{s}}{fk_{i}(\text{TVL}_{i}+\epsilon_{v}w_{i})}.$
Taking the power $\frac{1}{\epsilon_{v}-2}$ on both sides:
$\text{TVL}_{i}+w_{i}=\left(\frac{r_{s}}{fk_{i}(\text{TVL}_{i}+\epsilon_{v}w_{i})}\right)^{\frac{1}{\epsilon_{v}-2}}.$
Solving for $w_{i}$:
$w_{i}=\left(\frac{r_{s}}{fk_{i}(\text{TVL}_{i}+\epsilon_{v}w_{i})}\right)^{\frac{1}{\epsilon_{v}-2}}-\text{TVL}_{i}.$
Finally, the budget constraint must hold:
$W=w_{0}+\sum_{i=1}^{n}w_{i}.\squareforqed$